Circular No. 9034

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2009 F6 On Apr. 4, R. D. Matson informed the Central Bureau that he had

found a possible comet on ultraviolet SWAN images posted at the SOHO website, providing the rough positions below, and a search ephemeris was then posted at the Minor Planet Center’s ‘NEOCP’ webpage. Numerous CCD astrometrists and one visual observer have reported confirmation of the object as a comet, with selected positions provided below.

2009 UT  α_{2000}  δ_{2000}  Mag. Observer
Mar. 29  23^{h}03^{m}  +51.^{o}5  SWAN
31  23 16  +52.5  _{”}
Apr. 1  23 24  +53  _{”}
3  23 36  +53  _{”}
4  23 46  +54  _{”}
6.15  23 54.0  +54 41.^{′}5  8.2 Gonzalez
6.20771 23 54 17.^{s}90 +54 42 26′′.3 14.5 Hormuth
6.45823 23 56 18.14 +54 47 35.9 13.8 Guido
6.69785 23 58 12.89 +54 52 24.1 12.0 Kryachko
6.77483 23 58 50.08 +54 53 56.5 12.7 Ligustri
6.77683 23 58 51.85 +54 53 59.9  Ikari
6.80162 23 59 03.04 +54 54 28.3 13.5 Bittesini
6.80936 23 59 06.91 +54 54 36.7 11.3 Manca
6.84729 23 59 25.31 +54 55 21.3 13.0 Buzzi

J. J. Gonzalez (Burgos, Spain). Visual, 15-cm reflector; coma diameter 6′. F. Hormuth (Almería, Spain). 1.23-m reflector; 1′.5 coma with a well-

defined core, in twilight. E. Guido, G. Sostero, and P. Camilleri (Mayhill, NM, U.S.A.). Remotely

controlled 25-cm f/3.4 reflector; diffuse coma ∼ 2′.5 in diameter, with a sharp central condensation ∼ 20′′ in diameter.

T. Kryachko, S. Korotkiy, and B. Satovski (Zelenchukskaya). 30-cm reflec-

tor; bright 118′′ × 129′′ coma (major axis in p.a. 252^{o}) with condensation. R. Ligustri and P. Beltrame (Talmassons, Italy). 35-cm reflector. Coma

diameter ∼ 30′′. Y. Ikari (Moriyama, Shiga-ken, Japan). 26-cm f/7 reflector; 0′.7 coma. L. Bittesini, D. Di Maria, E. Pettarin, and F. Piani (Farra d’Isonzo, Italy).

30-cm f/4.5 reflector; coma diameter ∼ 50′′. F. Manca and A. Testa (Garbagnate Monastero, Italy). 50-cm reflector. L. Buzzi (Varese, Italy). 60-cm refl.; 1′.5 coma, strong central condensation.

2009 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green